# captionUnknown document class

### v2

@@ Abstract
Foundation models pre-trained using self-supervised learning have shown powerful transfer learning capabilities on various downstream tasks, including language understanding, text generation, and image recognition. The Earth observation (EO) field has produced several foundation models pre-trained directly on multispectral satellite imagery for applications like precision agriculture, wildfire and drought monitoring, and natural disaster response. However, few studies have investigated the ability of these models to generalize to new geographic locations, and potential concerns of geospatial bias—models trained on data-rich developed nations not transferring well to data-scarce developing nations—remain. We investigate the ability of popular EO foundation models to transfer to new geographic regions in the agricultural domain, where differences in farming practices and class imbalance make transfer learning particularly challenging. We first select five crop classification datasets across five continents, normalizing for dataset size and harmonizing classes to focus on four major cereal grains: maize, soybean, rice, and wheat. We then compare three popular foundation models, pre-trained on SSL4EO-S12, SatlasPretrain, and ImageNet, using in-distribution (ID) and out-of-distribution (OOD) evaluation. Experiments show that pre-trained weights designed explicitly for Sentinel-2, such as SSL4EO-S12, outperform general pre-trained weights like ImageNet. Furthermore, while only 100 labeled images are sufficient for achieving high overall accuracy, 900 images are required to achieve high average accuracy due to class imbalance. All harmonized datasets and experimental code are open-source and available for download.
@@ I Introduction
Transfer learning has been applied to remote sensing for agricultural monitoring and land cover mapping [16, 17], including detecting paddy rice in Spain and France and summer barley in the Netherlands [18], agricultural field boundary delineation [19], and geographic generalizability in tree detection [20]. Pre-trained foundation models have shown promising generalizability on downstream tasks with remote sensing data [21, 22, 23]. Remote sensing foundation models like SSL4EO-S12 [24] and SatlasPretrain [25] that are pre-trained on globally available data, such as Sentinel-1 and Sentinel-2, present a key opportunity for increasing the accuracy of global crop type mapping by improving the ability for models trained on downstream tasks to generalize across regions. However, transfer learning for crop type mapping in concert with such foundation models has not been studied at scale, in part because of the lack of a harmonized dataset of global crop type labels.
In this paper, we first present a harmonized global crop type mapping dataset that incorporates five regional datasets across five continents. We identify the available data sources with the highest label quality in each continent and pair their crop type labels with a uniform collection of cloud-free Sentinel-2 images captured during the peak of the growing season in each region. We focus on four major cereal grains and harmonize the class labels of each dataset around these categories. We then experiment with three popular pre-trained models and find that SSL4EO-S12 provides the best results. Further experiments with varying amounts of in-distribution and out-of-distribution data highlight the need for additional datasets in data-scarce regions like Africa and South America. All experiments are implemented using TorchGeo [26], all preprocessed datasets are available for download from Hugging Face [^1] https://huggingface.co/datasets/torchgeo/harmonized_global_crops, and all code is available on GitHub [^2] https://github.com/yichiac/crop-type-transfer-learning.

### v4

@@ Abstract
Foundation models pre-trained using self-supervised learning have shown powerful transfer learning capabilities on various downstream tasks, including language understanding, text generation, and image recognition. The Earth observation (EO) field has produced several foundation models pre-trained directly on multispectral satellite imagery for applications like precision agriculture, wildfire and drought monitoring, and natural disaster response. However, few studies have investigated the ability of these models to generalize to new geographic locations, and potential concerns of geospatial bias—models trained on data-rich developed nations not transferring well to data-scarce developing nations—remain. We evaluate three popular EO foundation models, SSL4EO-S12, SatlasPretrain, and ImageNet, on five crop classification datasets across five continents. Results show that pre-trained weights designed explicitly for Sentinel-2, such as SSL4EO-S12, outperform general pre-trained weights like ImageNet. While only 100 labeled images are sufficient for achieving high overall accuracy, 900 images are required to mitigate class imbalance and improve average accuracy.
@@ I Introduction
Transfer learning has been applied to remote sensing for agricultural monitoring and land cover mapping [16, 17], including detecting paddy rice in Spain and France and summer barley in the Netherlands [18], agricultural field boundary delineation [19], and geographic generalizability in tree detection [20]. Pre-trained foundation models have shown promising generalizability on downstream tasks with remote sensing data [21, 22, 23]. Remote sensing foundation models like SSL4EO-S12 [24] and SatlasPretrain [25] pre-trained on globally available data, such as Sentinel-1/2, present a key opportunity for increasing the accuracy of global crop type mapping by improving the ability for models trained on downstream tasks to generalize across regions. However, transfer learning for crop type mapping in concert with such foundation models has not been studied at scale, in part because of the lack of a harmonized dataset of global crop type labels.
@@
In this paper, we first present a harmonized global crop type mapping dataset that incorporates five regional datasets across five continents. We identify the available data sources with the highest label quality in each continent and pair their crop type labels with a uniform collection of cloud-free Sentinel-2 images captured during the peak of the growing season in each region. We focus on four major cereal grains and harmonize the class labels of each dataset around these categories. We then experiment with three popular pre-trained models and find that SSL4EO-S12 provides the best results. Further experiments with varying amounts of in-distribution and out-of-distribution data highlight the need for additional datasets in data-scarce regions like Africa and South America. All experiments are implemented using TorchGeo [31], all preprocessed datasets and code are available for download from Hugging Face [^1] https://huggingface.co/datasets/torchgeo/harmonized_global_crops and GitHub [^2] https://github.com/yichiac/crop-type-transfer-learning.